# Orders in Chaos: Enhancing Large-Scale MoE LLM Serving with Data Movement Forecasting

### v1

@@ 3.3.2. Expert Pair Co-activation Affinify
Figure 6. Co-activation probability heatmap of expert pair for (a) Deepseek and (b) Qwen, together with (c) statistical results across all layers, demonstrating that specific expert pairs exhibit high simultaneous activation probabilities.
3.3.3. System Insights from Spatial Relation.
@@
Hot experts vary across different tasks—experts for English queries differ from those for Chinese queries. Providing task-specific information enables proactive expert redistribution for improved performance. When serving predominantly English queries, English-relevant experts can be duplicated or reallocated beforehand. This task-to-expert mapping requires only one-time offline collection per model, making it practically viable.
4. Case Study: Future GPUs Architecture Design for MoE Models
@@
4.1. Trend of Future GPU Architecture
GPU vendors are increasingly adopting multi-chiplet architectures to overcome single-die performance limitations. As Moore’s Law approaches its limits (Lundstrom, 2003) and single-die size remains constrained by photomask dimensions (800-1,000 mm2), advanced packaging technologies like TSMC’s CoWoS (Hu et al., 2023), Samsung’s X-Cube (Shin et al., 2024), and Intel’s EMIB (Mahajan et al., 2016) enable multiple chiplets within a single package. Leading vendors have adopted such designs: AMD’s MI300 (Smith et al., 2024) integrates eight compute chiplets, NVIDIA’s Blackwell features two chiplets (NVIDIA, 2025a), and upcoming Rubin expects four (NVIDIA, 2025b).
This trend is evolving toward wafer-scale systems (Hou et al., 2017). TSMC’s System-on-Wafer (SoW) technology accommodates up to 24 compute dies and 96 HBM dies on a single wafer, exceeding 200,000 mm2 (Shih et al., 2025). Similar wafer-scale integration is also supported by Samsung I-Cube technology (Samsung, 2025). As shown in Figure 7(a), a typical wafer-scale multi-chiplet GPU consists of multiple units, each containing a GPU die and several HBM dies interconnected in a mesh topology. Such systems contain over 10 TB of HBM and PFLOPS-level computing power, supporting extremely large models and batch sizes.
The TSMC SoW technology shown in Figure 7(b) connects each GPU die to local HBM dies through local-silicon interconnects (LSI) (Hou et al., 2019), with adjacent GPU dies communicating via LSI vertically or XSR SerDes links horizontally. Although LSI and SerDes both provide terabit-level bandwidth, inter-GPU-die communication remains the primary bottleneck. Remote data access requires multiple hops across die-to-die links, resulting in high latency. Simultaneous remote HBM access by multiple dies creates bandwidth contention and traffic congestion, further degrading performance.
4.2. Challenges of serving MoE with future GPUs
@@
Simplistic Task Allocation. Current GPUs integrate a CPU in their SoC to serve as a command processor and allocate tasks to all SMs. However, the traditional command processors treat all SMs equally, ignoring their physical locations and data placement (LeBeane et al., 2018; Gutierrez et al., 2018). This causes tasks to be assigned to SMs far from the required data, generating excessive D2D traffic, and ignores MoE expert selection skewness, leading to poor utilization when most dies remain idle while others become overloaded.
@@
Figure 7. (a) Wafer-scale multi-chiplet GPU architecture with additional units highlighted in orange. (b) SoW (System-on-Wafer) technology structure. (c) Data format in the Global Command Processor for our proposed task distribution strategy.
4.3. Motivation and Insights
@@
4.4. Architecture Design
4.4.1. Architecture Overview
@@ 4.4.2. Key Data Structures
4.4.3. Workflow During Kernel Launch
@@
4.4.4. Dataflow for Remote Data Access
@@
4.4.5. Algorithm Design
Figure 8. Our proposed Task allocation algorithm (a) and data-driven predictor (b).
Table 2. Hardware Configurations
X-die Y-die
DRAM
BW
D2D
BW
DRAM
Cmpt Power
per Die (FP8)
Dojo (Talpes et al., 2023) 5 5 2 TB/s 1.5 TB/s 256 GB 1000 TFLOPS
TSMC-SoW (TSMC, 2025) 3 8 2 TB/s 1.5 TB/s 256 GB 1000 TFLOPS
5. Evaluation
Figure 9. Throughput of MoE layers (scaled to baseline).
Figure 10. Hop number reduction ratio (scaled to baseline).
5.1. Experiment Setup
Methodology: We developed a custom multi-chiplet GPU simulator in Python due to the lack of suitable existing tools. While cycle-accurate simulators like Gem5 (Binkert et al., 2011), gpgpusm (Bakhoda et al., 2009), and mgpusim (Sun et al., 2019) accurately model single GPUs, they are too slow for large-scale systems with over 20 GPUs serving batch sizes exceeding 15,000. ASTRI-sim (Won et al., 2023) supports multi-GPU systems but relies on real system traces and targets traditional LLM training, lacking MoE serving support.
Our trace-based simulator models key components including LLC, HBM, Compute Units, and D2D links across all dies while considering hardware resource contention. The simulator supports configurable die numbers, placement, and connectivity. To balance accuracy and efficiency, we simulate MoE models at expert slice granularity, with each expert comprising two slices.
Metric: We measure pure MoE layer performance during the decode stage as modern LLM serving systems show a trend toward fine-granularity disaggregation. Traditional LLMs benefit from separating prefill and decode stages across different machines, as demonstrated by DistServe (Zhong et al., 2024) and subsequent works (Wu et al., 2024; Qin et al., 2025). For MoE models, this disaggregation extends further. MegaScale-Infer (Zhu et al., 2025) separates attention and MoE operations onto different machines for optimal batch sizes. Following this trend, we focus on optimizing MoE operations during the decode stage.
Hardware Configuration: We evaluate on two multi-chiplet topologies: Tesla Dojo (Talpes et al., 2023, 2022) and TSMC SoW roadmap (TSMC, 2025). As shown in Table 2, Tesla Dojo features a 5×5 2D mesh while TSMC SoW employs an 8×3 2D mesh. We selected these configurations as they represent real-world deployment (Tesla Dojo) and near-future industry support (TSMC SoW). Each chiplet has 1,000 TFLOPS FP8 computing power, 512 GB HBM, 2 TB/s local HBM bandwidth, and 1.5 TB/s interconnection bandwidth between adjacent dies. We reserve 10% of DRAM for hardware management.
@@ 18 Function GenCandidateList(expert_idexpert\_id, disdis):
Models and Workloads: We conduct evaluations with real traces collected from Qwen3 and Deepseek V3. The traces are gathered from diverse datasets, including MMLU (Hendrycks et al., 2020), MMLU Pro (Wang et al., 2024), ChineseSimpleQA (He et al., 2024), and LiveCodeBench (Jain et al., 2024), comprising over 24,000 requests per model.
5.2. Throughput
@@
Comparison across configurations: Pred Only strategy achieves 40% performance improvement over baseline on average. It shows more substantial improvement on Qwen than Deepseek (68% vs. 14%) for two reasons. First, expert selection in Qwen is more predictable, making Pred Only more effective. Second, Qwen has smaller expert sizes, enabling us to cache more experts in local DRAM and reduce cross-unit communication.
5.3. Hop Reduction
@@
Pred Only reduces hop counts by 14%-78%, aligning with performance results in Figure 9. This indicates cross-unit communication is the primary bottleneck in baseline, and reducing hop counts proportionally improves performance.
@@
Figure 11. DRAM access breakdown for Qwen on TSMC with batch size 4096.
5.4. DRAM Access Breakdown
@@
5.5. Discussion
@@
This wafer-scale GPU architecture simply serves as a case study demonstrating the practical applicability of our proposed insights, which constitute the paper’s primary contribution. Our design implements only a subset of the proposed insights to validate their effectiveness. We follow Insight 3 for task allocation and leverage temporal relation insights (Insight 1 and part of Insight 2) to build a data-driven predictor. However, the remaining insights offer significant potential for further exploitation in wafer-scale architectures. For instance, Insight 2 could inform LLC prefetch strategies and enhance task allocation algorithms that consider LLC data placement. Spatial relation insights (Insight 4 and Insight 5) could optimize initial expert placement, which could be further refined by incorporating task information from Insight 6. Importantly, our insights extend far beyond this case study to benefit diverse future systems including multi-GPU clusters (Zhu et al., 2025; Zhang et al., 2025), CXL-based or CPU-based memory disaggregation (Cao et al., 2025a; Fang et al., 2025), Flash-based multi-hierarchy systems (Yu et al., 2024; Sheng et al., 2023), and other emerging architectures and systems.
6. Related Work
@@
MoE model behavior studies: Several MoE model tech reports (Muennighoff et al., 2024; Jiang et al., 2024; Dai et al., 2024; Sun et al., 2024) provide the MoE routing patterns as part of their evaluation. For example, the Mixtral report (Jiang et al., 2024) shows the temporal locality of expert selection by reporting the percentage of repetitive assignment. The OLMoE report (Muennighoff et al., 2024) shows the co-activation pattern and domain specialization among the experts. A blog post from SGLang (Team, [n. d.]) shows expert distribution statistics for the DeepSeekV3 model and the inherent imbalance in expert selection and similarity between prefill and decode. None of these studies provides a comprehensive profiling across multiple large-scale (¿200B) MoE models, nor do they present a data-movement-centric methodology to highlight the opportunities like this paper.
@@
Data movement optimization for MoE LLM inference Various prior work (Hwang et al., 2024; Li et al., 2023; Zhai et al., 2023; Gupta et al., 2024; Kamahori et al., 2025; Eliseev and Mazur, 2023; Du et al., 2024; Li et al., 2024; Skliar et al., 2024; Cao et al., 2025b; Yun et al., 2024; Zhang et al., 2025; Hwang et al., 2023) has focused on improving the efficiency of MoE LLM inference by reducing data movement. For example, Lina (Li et al., 2023) exploits the skewness in expert selection to dynamically schedule resources during inference and balance traffic across multiple GPUs. LYNX (Gupta et al., 2024) dynamically reduces active experts while preserving model accuracy. Pre-gate MoE (Hwang et al., 2024) uses pre-gating function to alleviate the dynamic nature of expert selection. Sida (Du et al., 2024) builds an offline hash function to predict expert usage to reduce data movement between CPU and GPU. MoE-Lightning (Cao et al., 2025b) leverages CPU-GPU pipeline and paged weights to improve resource utilization. Eliseev and Mazur (Eliseev and Mazur, 2023) exploit expert locality and leverage LRU caching to manage GPU and CPU memory. While they all aim to reduce data movement in MoE LLM inference of today’s systems, our case study shows that data-driven profiling across multiple models provides a solid base to optimize not only current but also future systems of any scale.
@@
Wafer-Scale and Chiplet Architectures As single chip scaling slows down, packaging technologies such as wafer-scale and chiplet packaging have become a promising path to further improve compute efficiency for future systems. Recent work in this area either focuses on the interconnect design (Rashidi et al., 2025; Yang et al., 2025b; Yu et al., 2025; Li and Wentzlaff, 2024; Chen et al., 2024; Feng et al., 2023a, b) of wafer-scale and chiplet systems, or general data placement for specific algorithms (He et al., 2025; Tan et al., 2021) and applications (Shao et al., 2019; Xu et al., 2025; Odema et al., 2024). In contrast, our work is the first to study the implications of serving large-scale MoE LLMs in a wafer-scale GPU and propose data-movement-centric HW/SW co-design optimizations.
7. Conclusion
@@
In this work, unlike previous MoE serving works that adopt system-centric approaches and develop deployment-specific strategies, we study MoE model serving from a model-focused perspective. We conducted comprehensive data-movement-centric profiling of large-scale MoE models (200B+) to extract system-agnostic insights. These insights reveal underlying patterns behind seemingly random data movement, providing actionable guidance for designing diverse future MoE serving systems. We validated several insights on wafer-scale GPU architectures and achieved significant performance improvements with minimal modifications.

### v2

@@ III-C2 Expert Pair Co-activation Affinify
Figure 8: Co-activation probability heatmap of expert pair for (a) Deepseek and (b) Qwen, together with (c) statistical results across all layers, demonstrating that specific expert pairs exhibit high simultaneous activation probabilities.
III-C3 System Insights from Spatial Relation.
@@
Hot experts vary by task and language. English queries, for instance, activate different expert subsets than Chinese queries. Providing task metadata during serving enables proactive expert placement: when workloads are predominantly English, systems can pre-duplicate or reassign English-relevant experts, reducing communication and balancing loads. This task-to-expert mapping requires only one-time offline profiling per model and can be reused throughout deployment, making the approach practical and efficient.
Figure 9: (a) Wafer-scale multi-chiplet GPU architecture with additional units highlighted in orange. (b) SoW (System-on-Wafer) technology structure. (c) Data format in the Global Command Processor for our proposed task distribution strategy.
IV Case Study: Future GPUs Architecture Design for MoE Models
@@
IV-A Trend of Future GPU Architecture
GPU vendors are increasingly adopting multi-chiplet architectures to overcome single-die performance limitations. As Moore’s Law approaches its limits [lundstrom2003moore] and single-die size remains constrained by photomask dimensions (800-1,000 mm2), advanced packaging technologies like TSMC’s CoWoS [hu2023cowos], Samsung’s X-Cube [xcube], and Intel’s EMIB [emib] enable multiple chiplets within a single package. Leading vendors have adopted such designs: AMD’s MI300 [mi300] integrates eight compute chiplets, NVIDIA’s Blackwell features two chiplets [blackwell], and upcoming Rubin expects four [rubin].
This trend is evolving toward wafer-scale systems [cowos2]. TSMC’s System-on-Wafer (SoW) technology accommodates up to 24 compute dies and 96 HBM dies on a single wafer, exceeding 200,000 mm2 [sow-x]. As shown in Figure 9(a), a typical wafer-scale multi-chiplet GPU consists of multiple units, each containing a GPU die and several HBM dies interconnected in a mesh topology. Such systems contain over 10 TB of HBM and PFLOPS-level computing power, supporting extremely large models and batch sizes.
The TSMC SoW technology shown in Figure 9(b) connects each GPU die to local HBM dies through local-silicon interconnects (LSI) [lsi], with adjacent GPU dies communicating via LSI vertically or XSR SerDes links horizontally. Although LSI and SerDes both provide terabit-level bandwidth, inter-GPU-die communication remains the primary bottleneck. Remote data access requires multiple hops across die-to-die links, resulting in high latency. Simultaneous remote HBM access by multiple dies creates bandwidth contention and traffic congestion, further degrading performance.
IV-B Challenges of serving MoE with future GPUs
@@
Simplistic Task Allocation. Current GPUs integrate a CPU in their SoC to serve as a command processor and allocate tasks to all SMs. However, the traditional command processors treat all SMs equally, ignoring their physical locations and data placement [cp1, cp2]. This oblivious task-to-SM assignment generates excessive D2D traffic and ignores MoE expert selection skewness, leading to poor utilization when most dies remain idle while others become overloaded.
@@
IV-C Motivation and Insights
@@
IV-D Architecture Design
IV-D1 Architecture Overview
@@ IV-D2 Key Data Structures
IV-D3 Workflow During Kernel Launch
@@
Figure 10: Proposed task allocation algorithm and data-driven predictor.
IV-D4 Dataflow for Remote Data Access
@@ 18 Function GenCandidateList(expert_idexpert\_id, disdis):
IV-D5 Algorithm Design
V Evaluation
Figure 11: Throughput of MoE layers (Top) and hop number reduction ratio (Bottom). All figures are scaled to baseline.
V-A Experiment Setup
Methodology: We conduct experiments using trace-driven simulation on a validated simulator. Expert selection traces are collected by deploying SGLang [zheng2024sglang] on an 8×H100 DGX server and an 8×H200 AWS instance.
We developed a custom multi-chiplet GPU simulator in Python since existing tools are inadequate for our needs. Cycle-accurate simulators like Gem5 [gem5], gpgpusim [gpgpusim], and mgpusim [mgpusim] accurately model single GPUs but are too slow for large-scale systems with 20+ GPUs serving batch sizes exceeding 15,000. Event-driven ASTRA-sim [astra] supports multi-GPU systems but lacks MoE serving support and detailed single-GPU microarchitecture modeling. To ensure accuracy, we validated our simulator against real data from an 8×H100 DGX server, as shown in Sec. V-B.
Figure 12: Simulator validation with real data generated from 8xH100 DGX, including both MoE GEMM (Top) and P2P data transfer (Bottom) test cases.
Our simulator models multi-chiplet GPU components, including LLC, HBM, compute units, and D2D links across all dies. A central resource manager simulates contention and congestion by dividing each expert execution into multiple events and managing simultaneous execution of thousands of events. The simulator supports configurable die numbers, placement, and connectivity. To balance accuracy and efficiency, we simulate MoE models at expert slice granularity, with each expert comprising two slices.
TABLE I: Hardware Configurations
X-die Y-die
DRAM
BW
D2D
BW
DRAM
Cmpt Power
per Die (FP8)
Dojo 5 5 2 TB/s 1.5 TB/s 80GB 1000 TFLOPS
TSMC-SoW 3 8 2 TB/s 1.5 TB/s 80GB 1000 TFLOPS
Dojo-Enhanced 5 5 8 TB/s 2 TB/s 180GB 4500 TFLOPS
Other Params
LLC hit latency: 100ns, LLC miss penalty: 110ns,
LLC write latency: 30ns, LLC size: 64 MB
D2D link latency: 200ns, Routing Alg: XY routing,
Command and address size for each remote request: 16B
Loca HBM access latency: 300 ns
Metric: We measure pure MoE layer performance during the decode stage as modern LLM serving systems show a trend toward fine-granularity disaggregation. Traditional LLMs benefit from separating prefill and decode stages across different machines, as demonstrated by DistServe [zhong2024distserve] and subsequent works [loongserve, qin2025mooncake]. For MoE models, this disaggregation extends further. MegaScale-Infer [zhu2025megascale-infer] separates attention and MoE operations onto different machines for optimal batch sizes. Following this trend, we focus on optimizing MoE operations during the decode stage.
[…]
For both the Dojo and TSMC SoW configurations, each chiplet is A100-like, providing 1,000 TFLOPS FP8 compute, 80GB HBM, 2TB/s local HBM bandwidth, and 1.5 TB/s inter-die bandwidth to adjacent chiplets. We also include an extended experiment in Sec. V-F with a Dojo-Enhanced configuration, where each die is B300-like to reflect an anticipated hardware performance trend in the future. We reserve 10% of DRAM for system and hardware management.
@@ V-A Experiment Setup
Models and Workloads: We conduct evaluations with real traces collected from Qwen3 and Deepseek V3. The traces are gathered from diverse datasets, including MMLU [mmlu], MMLU Pro [mmlu_pro], ChineseSimpleQA [he2024chinese], and LiveCodeBench [jain2024livecodebench], comprising over 24,000 requests per model.
V-B Validation of Simulator
We validate our simulator using real measurements from an 8×H100 DGX server. We evaluate both single-GPU execution and two-GPU peer-to-peer (P2P) communication. For single-GPU execution, we benchmark GEMM operations for experts in the MoE layer across varying batch sizes for both DeepSeek and Qwen. For P2P communication, we measure data migration between two GPUs over payload sizes ranging from 4 KB to 4 GB. To ensure simulation fidelity, we calibrate key parameters to fit the measured data. As shown in Figure 12, the simulator’s error remains within 5% for all test cases.
V-C Throughput
@@
V-D Hop Reduction
@@
[…]
V-E DRAM Access Breakdown
@@
Figure 13: DRAM access breakdown for Qwen3 on TSMC-SoW Configuration with batch size 4096.
Figure 14: Host CPU implementation overhead under varying models and batch sizes.
V-F Comparison with Host CPU-Based Implementation
Our task allocation algorithm runs on a new GPU command processor, but it could, in principle, be executed on the host CPU with a higher overhead. As shown in Figure 14, we evaluate both the Dojo and Dojo-Enhanced configurations. In Dojo, the overhead of host-CPU allocation is 5.2%–6.4% for DeepSeek V3 and 11.1%–14.2% for Qwen3. In Dojo-Enhanced, the overhead rises to 19.3%–23.8% for DeepSeek V3 and 42.0%–51.6% for Qwen3.
DeepSeek vs Qwen: Qwen3 incurs higher overhead than DeepSeek V3 due to CPU–GPU data transfers over PCIe, which occur once per MoE layer. The CPU needs the Expert Distribution Table from the GPU to run the allocator, and the allocation results must be sent back to the GPU before kernel execution. Qwen3 has (i) more MoE layers (94 vs. 58), increasing transfer frequency, and (ii) smaller per-layer compute, which amplifies the relative cost of transfers.
Dojo vs Dojo-Enhanced: Dojo-Enhanced shows over 3.7× higher overhead than Dojo because its GPU dies are significantly faster, making fixed PCIe transfer costs dominate more. As GPU performance outpaces interconnect bandwidth, implementing the allocator in the GPU command processor becomes increasingly necessary to sustain performance.
V-G Discussion
@@
This wafer-scale GPU architecture serves as a case study demonstrating the practical applicability of our proposed insights, which constitute the paper’s primary contribution. Our design implements only a subset of the proposed insights to validate their effectiveness. We follow Insight 3 for task allocation and leverage temporal relation insights (Insight 1 and part of Insight 2) to build a data-driven predictor. However, the remaining insights offer significant potential for further exploitation in wafer-scale architectures. For instance, Insight 2 could inform LLC prefetch strategies and enhance task allocation algorithms that consider LLC data placement. Spatial relation insights (Insight 4 and Insight 5) could optimize initial expert placement, which could be further refined by incorporating task information from Insight 6. Importantly, our insights extend far beyond this case study to benefit diverse future systems, including multi-GPU clusters [zhu2025megascale-infer, zhang2025comet], CXL-based or CPU-based memory disaggregation [moelightning, fang2025klotski], Flash-based multi-hierarchy systems [cambricon-llm, sheng2023flexgen], and other emerging architectures and systems.
VI Related Work
@@
MoE model behavior studies: Several MoE model tech reports [muennighoff2024olmoe, jiang2024mixtral, dai2024deepseekmoe, sun2024hunyuan] provide the MoE routing patterns as part of their evaluation. For example, the Mixtral report [jiang2024mixtral] shows the temporal locality of expert selection by reporting the percentage of repetitive assignment. The OLMoE report [muennighoff2024olmoe] shows the co-activation pattern and domain specialization among the experts. A blog post from SGLang [sglang-deepseek-blogpost] shows expert distribution statistics for the DeepSeekV3 model and the inherent imbalance in expert selection and similarity between prefill and decode. None of these studies provides a comprehensive profiling across multiple large-scale (>>200B) MoE models, nor do they present a data-movement-centric methodology to highlight the opportunities like this paper.
@@
Data movement optimization for MoE LLM inference Various prior work [hwang2024pre, li2023accelerating, zhai2023smartmoe, gupta2024lynx, kamahori2025fiddler, eliseev2023fast, du2024sida, li2024optimizing, skliar2024mixture, moelightning, yun2024duplex, zhang2025comet, hwang2023tutel] have focused on improving the efficiency of MoE LLM inference by reducing data movement. For example, Lina [li2023accelerating] exploits the skewness in expert selection to dynamically schedule resources during inference and balance traffic across multiple GPUs. LYNX [gupta2024lynx] dynamically reduces active experts while preserving model accuracy. Pre-gate MoE [hwang2024pre] uses a pre-gating function to alleviate the dynamic nature of expert selection. Sida [du2024sida] builds an offline hash function to predict expert usage to reduce data movement between CPU and GPU. MoE-Lightning [moelightning] leverages a CPU-GPU pipeline and paged weights to improve resource utilization. Eliseev and Mazur [eliseev2023fast] exploit expert locality and leverage LRU caching to manage GPU and CPU memory. This work aims to reduce data movement in MoE LLM inference of today’s systems, and our case study shows that data-driven profiling across multiple models provides a solid base to optimize not only current but future systems of any scale.
@@
Wafer-Scale and Chiplet Architectures As single chip scaling slows down, packaging technologies such as wafer-scale and chiplet packaging have become a promising path to further improve compute efficiency for future systems. Recent work in this area either focuses on the interconnect design [rashidi2025fred, yang2025pd, yu2025cramming, li2024lucie, chen2024waferscale, feng2023heterogeneous, feng2023scalable] of wafer-scale and chiplet systems, or general data placement for specific algorithms [he2025waferlm, tan2021nn] and applications [shao2019simba, xu2025wsc, odema2024scar]. In contrast, our work is the first to study the implications of serving large-scale MoE LLMs in a wafer-scale GPU and propose data-movement-centric HW/SW co-design optimizations.
VII Conclusion
@@
In this work, unlike previous MoE serving works that adopt system-centric approaches and develop deployment-specific strategies, we study MoE model serving from a model-focused perspective We conducted comprehensive data-movement-centric profiling of the latest large-scale MoE models (200B-1000B) to extract system-agnostic insights. These insights reveal underlying patterns behind seemingly random data movement, providing actionable guidance for designing diverse future MoE serving systems. We validate our insights on a wafer-scale GPU architecture and achieve significant performance improvements with minimal modifications.